The voltage of the following cell at 25 ̊C is + 0.65V. The standard reduction potential for Zn is -0.76 V.

Zn | Zn2+ (0.60M) || H+ (aq.), H2 (g,1.0 atm) | Pt

What is the pH of the solution at the cathode?
Answer #1

Genen: Ecell = 0.65V Zn 2t+ae (09) -> Zn; E cell = -0.76V - We know for hydrogens electrode 2 Ht the Arhi E cell = 0 + Zn 12/pH = -logent) - log (0.12) ..ph 0.921 Are CS Scanned with CamScanner